OpenStudy (anonymous):

Determine if (3/4' - 2) is a solution to -4x-7y =11 . Please show all of your work.

OpenStudy (anonymous):

substitute into the equation x=3/4, y=-2 -4(3/4)-7*(-2)=11 -3+14=11 so it's a solution

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes